A local retailer recently collected data on customers that frequent the store. The data are presented in the table below

Items Net Sales Gender Age
1 59.25 Male 29
1 153.6 Female 33
1 33.75 Female 29
5 150.6 Female 25
2 81 Female 31
1 66.75 Female 41
2 117 Female 27
1 33.75 Female 37
2 84.78 Female 43
1 66.75 Female 33
1 44.25 Female 45
1 47.4 Female 37
9 240.6 Female 37
2 96.75 Female 43
1 74.25 Male 21
2 107.1 Male 33
3 141 Female 19
3 81.75 Female 37
2 57.75 Female 29
6 67.2 Female 53
1 47.4 Female 25
4 106.23 Female 35
7 399 Female 47
2 111 Female 39
2 59.25 Male 45
1 45.03 Female 57
1 66.75 Female 51
5 289.2 Female 39
3 106.8 Female 29
1 27 Female 67

a. Develop a frequency distribution for the number of Items purchased. What is the probability that a customer will purchase 3 or fewer items? b. develop a frequency distribution and histogram for Net Sales using a class width of 25 (1-25, 26-50...). c. Use a pie chart to display the percentages of customers who are male or female. d. for the age data, determine the following: (1) mean: (2) Median (3) Mode (4) Interquartile range (5) 94nd percentile
Answer #1

(a)

Items Frequency
1 13
2 8
3 3
4 1
5 2
6 1
7 1
8 0
9 1


Here, total frequency = 30. Hence, required probability = (13 + 8 + 3) = 24/30 = 0.80.

(c) The pie chart is given below:


(b)

Net Sales Frequency
1-25 0
26-50 7
51-75 8
76-100 4
101-125 5
126-150 2
151-175 1
176-200 0
201-225 0
226-250 1
251-275 0
276-300 1
301-325 0
326-350 0
351-375 0
376-400 1




(d) Mean = 37.2, Median = 37, Mode = 29, IQR = Q3 - Q1 = 43 - 29 = 14, 94th percentile = 54.04.